Back to main Archive News page.
Thu 28 Sep 2006 - Downtime for part of Exchange Email system
12:00-14:00 We are investigating some potential hardware failures on the EXCHANGE-BE4 system, and our vendor support has advised us that some firmware updates are needed. The EXCHANGE-BE4 system was unavailable between 1200 and 1400 on Thursday September 28. Users whose mailboxes are on other servers were not be affected.
Wed 27 Sep 2006 - LUVLE and My Modules maintenance
14:00-15:00 LUVLE and My Modules was unavailable from 2pm today, 27th September 2006 for approximately 1 hour to allow for essential server maintenance work to be carried out.
Tue 26 Sep 2006 - ResNet Upgrade
07:00-09:00 During the network at risk period on the 26th September the router feeding ResNet was upgraded. This should provide improved performance to all users.
Mon 25 Sep 2006 - ResNet systems downtime
As part of work being done to prepare ResNet for the new academic year, there was a scheduled outage of around an hour on the ResNet system on Monday morning. All users will now be required to revalidate for the new academic year.
Fri 22 Sep 2006 - All systems ran without major problems
Thu 21 Sep 2006 - All systems ran without major problems
Wed 20 Sep 2006 - LUVLE and My Modules maintenance
13:30-14:30 LUVLE and My Modules was unavailable from 1:30pm to allow for essential server maintenance work to be carried out.
Tue 19 Sep 2006 - Network Maintenance in Fylde
07:00-09:00 During the at risk period on the 19th September the core router was upgraded in Fylde.
Mon 18 Sep 2006 - All systems ran without major problems
Sat 16 Sep 2006 - Network Problem in the Library
Due to a hardware failure the Library had no access to network resources until noon.
A planned hardware upgrade was bought forward due to this failure and full service was restored.
Fri 15 Sep 2006 - All systems ran without major problems
Thu 14 Sep 2006 - Problem with Central Filestore (H: drive), Power Loss to Great Hall and Round House, Questionmark Perception Outage
10:00 There is an unknown problem with central-files that is currently being investigated by ISS. More information will be provided when available.
11:00 Due to a power supply problem users in the Great Hall & Round House may experience network connectivity problems. This is outside ISS control, and University Estates have been informed.
The Questionmark Perception service was unavailable all day on Thursday 14th September for upgrading and re-siting the server. For details please contact Tim Ellis in CELT (email@example.com, ex 94835).
Wed 13 Sep 2006 - Network and LUVLE/MyModules Maintenance
There were brief interruptions to availability of central servers between 2pm and 4pm on Wednesday 13th September as part of the work to investigate intermittent core switch problems.
14:00-15:00 LUVLE and My Modules was unavailable for approximately 1hr to allow for essential server maintenance work to be carried out.
Tue 12 Sep 2006 - H: Drive Problem, & Network Maintenance
An issue relating to the server HOMES3 which serves about one third of users H: drives is currently being investigated. ISS hope to have service resumed shortly. Update (09:15) - HOMES3 is currently being repaired; we expect to return it to service by 11:00 this morning. Update 09:56 - HOMES3 is now back in service.
During the Network at Risk period on Tuesday there was a brief loss of the network across campus whilst essential changes were made. ISS are continuing to investigate intermittent core switch problems, with the assistance of the switch manufacturer
Mon 11 Sep 2006 - Network Problems across campus
09:50 ISS are continuing to investigate issues with a core network switch, which are resulting in periodic network outages.
Sat 09 Sep 2006 - Network Maintenance
14:00-14:30: Minor network outages were experienced while essential diagnostics were performed on a core switch.
Fri 08 Sep 2006 - Network Problems, and LUVLE/MtModules Maintenance
There were problems with a core switch affecting networking on campus.
14:00-15:00: LUVLE and My Modules were not available to allow for essential server maintenance work to be carried out.
Thu 07 Sep 2006 - All systems ran without major problems
Wed 06 Sep 2006 - ResNet Maintenance
During the Systems At Risk period, updates were applied to the ResNet management system. This resulted in short outages while some servers were rebooted
Tue 05 Sep 2006 - Power Outage in IAS
There was an unexpected power outage in the IAS building for part of the morning. Faculty servers were switched off overnight from 17:30 to allow Estates to perform remedial work.